# One Patch to Caption Them All: A Unified Zero-Shot Captioning Framework

### v2

@@ Abstract
Zero-shot captioners are recently proposed models that utilize common-space vision-language representations to caption images without relying on paired image-text data. To caption an image, they proceed by textually decoding a text-aligned image feature, but they limit their scope to global representations and whole-image captions. We present Patch-ioner, a unified framework for zero-shot captioning that shifts from an image-centric to a patch-centric paradigm, enabling the captioning of arbitrary regions without the need of region-level supervision. Instead of relying on global image representations, we treat individual patches as atomic captioning units and aggregate them to describe arbitrary regions, from single patches to non-contiguous areas and entire images. We analyze the key ingredients that enable current latent captioners to work in our novel proposed framework. Experiments demonstrate that backbones producing meaningful, dense visual features, such as DINO, are key to achieving state-of-the-art performance in multiple region-based captioning tasks. Compared to other baselines and state-of-the-art competitors, our models achieve better performance on zero-shot dense, region-set, and a newly introduced trace captioning task, highlighting the effectiveness of patch-wise semantic representations for scalable caption generation. Project page at paciosoft.com/Patch-ioner/.
@@ 1 Introduction
Image captioning is one of the most representative tasks in vision-language understanding and has reached incredible accuracy thanks to the availability of pre-trained vision-language backbones and large paired image-text datasets. In its basic formulation, a captioning model takes a full image as input and autonomously decides which elements must be described and up to what degree. To enable user guidance and produce more targeted descriptions, some previous works proposed region-level captioning methods (Johnson et al., 2016; Cornia et al., 2019), which take as an additional input a spatial indication — e.g., bounding boxes — specifying which image regions have to be described and, possibly, in which order.
@@
In this paper, we propose a perspective shift that enables us to perform region-level captioning with arbitrary spatial granularity — from a single image patch up to the entire image — in a zero-shot fashion, i.e., without requiring any form of image-level or region-level supervision. Specifically, instead of relying on the idea that the subject of a captioning method is the image — then potentially conditioned on a set of sub-regions — we instead build on two straightforward yet powerful ideas: i) the simplest element that we could caption is a patch, the atomic element of an image representation in modern architectures based on vision transformers (Dosovitskiy et al., 2021), and ii) we can easily aggregate multiple patch representations to produce descriptions for arbitrarily large — and also potentially not contiguous — image regions. We present Patch-ioner, a zero-shot captioning framework implementing these ideas. Our formulation offers maximum flexibility in zero-shot captioning tasks, producing models that effortlessly generate captions for various aggregations of image patches, ranging from individual patches to larger image regions, up to providing a caption for the entire image.
@@
In the last years, large pre-trained vision-language foundation models like CLIP (Radford et al., 2021; Jia et al., 2021; Li et al., 2023a) solved many downstream tasks in zero-shot or even training-free configurations. In particular, contrastively learned vision-language representations enabled impressive results in zero-shot settings in image classification (Radford et al., 2021; Zhai et al., 2022), open-vocabulary detection (Zhou et al., 2022; Minderer et al., 2023), and segmentation (Ghiasi et al., 2022; Liang et al., 2023), or text-image retrieval (Kordopatis-Zilos et al., 2025). Image captioning, however, cannot directly employ CLIP machinery at inference time to generate text, given that CLIP is inherently a discriminative — and not a generative — approach. Only recently, image captioning models became zero-shot by decoupling image encoding — where pre-trained discriminative models like CLIP are used to create proper image and text representations — from the actual generative module. This is the case for models like  Nukrai et al. (2022); Li et al. (2023b); Gu et al. (2023); Zeng et al. (2024); Fei et al. (2023); Yan et al. (2025); Zeng et al. (2025); Tewel et al. (2022); Su et al. (2022), which i) employ CLIP to leverage a shared vision-language semantic space, and ii) train a text decoder on solely text samples to recover the text back from the CLIP textual feature. This requires nothing more than a pre-trained contrastive model and a large set of sole text samples to craft a powerful captioner.
@@
In this paper, we show that our framework can generalize this core idea and that, under this formulation, many zero-shot captioners, paired with the right components, can be easily restructured to perform zero-shot region-based captioning. Therefore, we identify and study in detail the most critical components of this novel zero-shot region-based captioning framework. Particularly, we focus our attention on the pre-trained vision-language contrastive backbone, which should be able, unlike CLIP, to create meaningful patch representations. To this aim, we largely explore DINO-based  (Caron et al., 2021; Oquab et al., 2024) variants, having better localized capabilities than CLIP. We further expand the study of our framework, addressing multiple modality-gap mitigation strategies that help the text decoder to be trained only on text without having access to paired image-text features, as well as studying different patch aggregation methods.
@@
By studying existing components and employing vision backbones able to output patch-level meaningful representations like DINO, we show that we can enable many zero-shot captioners to reach state-of-the-art or comparable results in many zero-shot captioning task variants requiring captioning sub-parts of the entire image — dense captioning (Johnson et al., 2016), region-set captioning (Cornia et al., 2019), up to the standard image captioning (Tewel et al., 2022) where the region to caption extends over the entire image. To better showcase the effectiveness of our framework in extreme patch-based captioning scenarios, we also introduce the zero-shot trace captioning task, which requires generating a caption for a region within an image specified by a mouse trace.
@@
To summarize, we propose the following contributions: a) We reformulate captioning by shifting perspective from the image-to-caption approach to a patch-to-caption one, unifying local and global tasks in Patch-ioner, a single framework which does not require region-level supervision, b) we repurpose existing models to work within this novel framework, by analyzing the role of the key components, with a special attention to the vision backbone, c) we show the performance of these models on four zero-shot captioning tasks that span different region granularity, from captioning few patches to the whole image, showing the effectiveness of the proposed perspective shift proposed by our framework despite its overall simplicity.
@@
Figure 1: Patch-ioner: a patch-centric framework for unified zero-shot captioning. A. Overview of our framework. First, we extract language-aligned dense patch embeddings from the image using a VLM. Given a region, we select the underlying patches and aggregate their features to obtain a region representation. Finally, we obtain the region caption by applying a zero-shot text decoder, that is a) conditioned on the latent region representation, b) trained on text-only data, and c) equipped with a mechanism to handle the modality gap present in vision-language common spaces. This enables regional captioning without requiring region-level supervision. B. By aggregating patch-level features from arbitrary image regions, we can flexibly handle multiple captioning tasks across spatial granularities in a unique model.
@@ 2 Related Work
Language-aligned Dense Image Representations are crucial for our goal of captioning at patch level. Vision-language models (VLM) like CLIP (Radford et al., 2021) introduced a powerful approach to learning global modality representations in a shared space via contrastive learning, paving the way to solve several downstream tasks, including captioning (Mokady et al., 2021; Cornia et al., 2024). However, in zero-shot settings, CLIP-like representations are known to struggle with dense tasks due to misalignment between local visual patches and fine-grained semantics (Zhong et al., 2022; Ranasinghe et al., 2023; Bica et al., 2024). On the other hand, visual-only self-supervised models (SSM) like DINO (Caron et al., 2021; Oquab et al., 2024) excel in local semantic modeling but lack a bridge with language. Recent works like SILC (Naeem et al., 2024) and DINO.txt (Jose et al., 2024) attempt to get the best of both worlds by combining DINO- and CLIP-like training objectives, aiming to obtain language-aligned dense representations. INViTE (Chen et al., 2024) modifies CLIP’s visual encoder by zeroing the attention weights from each patch to all the others in the last layers, leading to more semantic patch features. DenseCLIP (Rao et al., 2022) and RegionCLIP (Zhong et al., 2022) extend CLIP with additional region-level supervision. RegionCLIP leverages a region-proposal network to construct region-text pairs from image-text datasets, while DenseCLIP introduces a pixel-to-text matching loss to strengthen the alignment between local regions and textual concepts. Other methods instead exploit already existing VLMs and SSMs to get the same properties with minimal or no training: Talk2DINO (Barsellotti et al., 2024) connects language to the DINOv2 space by mapping CLIP textual representations to DINOv2 patches. ProxyCLIP (Lan et al., 2024) instead leverages DINO’s attention maps to improve the local properties of the CLIP visual embeddings of patches.
@@
Zero-shot Image Captioning methods rely mostly on global CLIP representations to guide text generation. Early-guided decoding methods take CLIP visual features as input and introduce adaptation techniques to reduce the visual-textual modality gap (Liang et al., 2022). DeCap (Li et al., 2023b) projects CLIP visual features into a more text-aligned space using a memory of texts as basis, while CapDec (Nukrai et al., 2022) and CLOSE (Gu et al., 2023) inject noise during text-only training to enable decoding also from the CLIP visual space. To improve the generation ViECap (Fei et al., 2023), MeaCap (Zeng et al., 2024), MERCap (Zeng et al., 2025) and EntroCap (Yan et al., 2025) leverage external knowledge to condition the decoding together with the CLIP image representation. Late-guided decoding methods instead use CLIP as a scoring or optimization signal rather than direct input. ZeroCap (Tewel et al., 2022) leverages CLIP gradients to steer the cached context during text generation, while MAGIC (Su et al., 2022) optimizes token selection based on CLIP similarity scores. However, all the above approaches rely on global representations, which are not well-suited for capturing localized semantic details, making them suboptimal for patch-level or region-level captioning in zero-shot settings.
@@
Region-level Captioning comprises several tasks in which models are asked to produce natural language descriptions based on sub-parts of an image. They pose additional challenges as naively captioning the cropped regions or feature maps often induces a loss of the global context of the image and, thus, misinterpretation of the region. For this reason, zero-shot solutions to this family of problems are still underexplored. For controllable captioning (Cornia et al., 2019) — the generation of an image caption controlled by a set or sequence of regions — and dense captioning (Johnson et al., 2016) — the localization and captioning of salient regions of an image — state-of-the-art solutions like CAG-Net (Yin et al., 2019), GRiT (Wu et al., 2024), ControlCap (Zhao et al., 2024), and FlexCap (Dwibedi et al., 2025) provide good performance but need supervision with ground-truth boxes. Recent works (Guo et al., 2024), (Hua et al., 2025) moved towards the direction of arbitrary regions captioning exploiting region-level supervision, and the Localized Narratives dataset Pont-Tuset et al. (2020a) — comprising images, timed captions, and timed mouse tracks — provide the ingredients for evaluating captioning also at track- or patch-level. We propose a unique framework to tackle captioning at various granularities, from image- to patch-level, in a zero-shot setting.
@@ 3.1 Motivation and Formulation
Traditional regional captioning (Johnson et al., 2016; Dwibedi et al., 2025; Zhao et al., 2024) follows an early injection of region specification in the model. Formally, given an image $I$ and a region $R$, the region caption $t$ is modeled as $t=\mathcal{D}(I,R)$. Such models require region-caption annotations and often use dedicated models or losses per captioning task or granularity. A step forward can be moved by introducing a formulation which disentangles image encoding and postpones region specification, defining $t=\mathcal{D}(\psi(I),R)$, where $\psi(I)$ provides a visual representation of the image independent of the region $R$, and $\mathcal{D}$ performs late region selection and text decoding. In order to avoid training the whole pipeline using region-level labels, we further decompose the decoder module $\mathcal{D}(\cdot)$ into two distinct modules: a parameter-free fixed aggregation $\operatorname{agg}_{R}$ of patch-level representations, and an actual text decoder $\phi(\cdot)$ not directly conditioned on regions $R$, so that $t=\phi(\operatorname{agg}_{R}(\psi(I),R))$. We will detail these two factorized components in the following paragraphs.
@@
We train a text decoder $\phi:\mathbb{R}^{D}\to\mathcal{T}$ using a prefix language modeling approach, where the decoder reconstructs a caption $t\in\mathcal{T}$ from its text embedding $\psi_{\text{t}}(t)$. If $\psi_{\text{t}}(t)$ is aligned to the visual encoder $\psi_{v}(I)$, we could directly decode the visual embeddings using a text-only decoder trained on text-only data. In such a case, we can finally decode the region representation into a natural language caption $t=\phi(\mathbf{v}_{R})$. However, the assumption that $\phi(\cdot)$ can digest features from $\psi_{v}(I)$ while being trained to reconstruct features from $\psi_{\text{t}}(t)$ is often too optimistic due to the prominent modality gap (Liang et al., 2022). In fact, text and image representations, despite being in the same multimodal space, occupy different, separated subspaces. To obtain a decoder applicable to visual embeddings, we experimented with two mitigation strategies for this gap. The first strategy, following Li et al. (2023b), introduces a projection step at inference that maps visual features into the text subspace using a memory of text embeddings. The second, inspired by Gu et al. (2023) and Nukrai et al. (2022), trains the decoder under input perturbations, encouraging robustness so that it can directly process visual embeddings. We analyze the effects of the modality gap mitigation strategies in SM§9.
Overall, our formulation offers three main advantages: a) it is zero-shot by design, in the sense that it only requires image-level textual descriptions to be trained and does not require paired image-text samples to train the text decoder, b) any region (e.g., whole image, box, mask, free-form trace, single point) is addressed identically, supporting a modular, general-purpose regional captioning pipeline, and c) our method requires only a single forward pass of the vision backbone to extract patch features for an entire image, which can then be reused to caption multiple regions without rerunning the full pipeline for each one.

### v3

@@ Abstract
Zero-shot captioners are recently proposed models that utilize common-space vision-language representations to caption images without relying on paired image-text data. To caption an image, they proceed by textually decoding a text-aligned image feature, but they limit their scope to global representations and whole-image captions. We present a unified framework for zero-shot captioning that shifts from an image-centric to a patch-centric paradigm, enabling the captioning of arbitrary regions without the need of region-level supervision. Instead of relying on global image representations, we treat individual patches as atomic captioning units and aggregate them to describe arbitrary regions, from single patches to non-contiguous areas and entire images. We analyze the key ingredients that enable current latent captioners to work in our novel proposed framework. Experiments demonstrate that backbones producing meaningful, dense visual features, such as DINO, are key to achieving state-of-the-art performance in multiple region-based captioning tasks. Compared to other baselines and state-of-the-art competitors, our models achieve better performance on zero-shot dense captioning and region-set captioning. We also introduce a new trace captioning task that further demonstrates the effectiveness of patch-wise semantic representations for flexible caption generation. Data and code are available at https://paciosoft.com/Patch-ioner/.
@@ 1 Introduction
Image captioning is one of the most representative tasks in vision-language understanding and has reached outstanding accuracy thanks to the availability of pre-trained vision-language backbones and large paired image-text datasets. In its basic formulation, a captioning model takes a full image as input and autonomously decides which elements must be described and up to what degree. To enable user guidance and produce more targeted descriptions, some previous works proposed region-level captioning methods [23, 10], which take as an additional input a spatial indication (e.g., bounding boxes) specifying which image regions have to be described and, possibly, in which order.
@@
In this paper, we propose a perspective shift that enables us to perform region-level captioning with arbitrary spatial granularity — from a single image patch up to the entire image — without requiring any form of region-level supervision or paired text-region data. Specifically, instead of sticking with the classic setup where the subject of a captioning method is the image — then potentially conditioned on a set of sub-regions — we instead build on two straightforward yet powerful ideas: i) the simplest element that we could caption is a patch, the atomic element of an image representation in modern architectures based on vision transformers [12], and ii) we can easily aggregate multiple patch representations from frozen vision-language backbones to produce descriptions for arbitrarily large — and also potentially not contiguous — image regions.
We present a regional captioning framework which implements these ideas through a zero-shot regional captioning setup: we build on the assumptions that paired text-region data is not available at training time, and the models can be solely trained on text data, assumed to be available in large quantities. Our formulation offers maximum flexibility in zero-shot regional captioning tasks, producing models that effortlessly generate captions for various aggregations of image patches, ranging from individual patches to larger image regions, up to providing a caption for the entire image.
@@
In the last years, vision-language foundation models like CLIP [50, 22, 32] solved many downstream tasks in zero-shot or even training-free configurations. In particular, contrastively learned vision-language representations enabled impressive results in zero-shot settings in image classification [50, 69], open-vocabulary detection [72, 39], and segmentation [16, 35], or text-image retrieval [27]. Image captioning, however, cannot directly employ CLIP machinery at inference time to generate text, given that CLIP is inherently a discriminative — and not a generative — approach. Only recently, image captioning models became zero-shot by decoupling image encoding — where pre-trained vision-language models like CLIP are used to create proper image and text representations — from the actual generative module. This is the case for models like  [43, 33, 18, 68, 15, 63, 67, 59, 56], which i) employ CLIP to leverage a shared vision-language semantic space, and ii) train a text decoder on solely text samples to recover the text back from the CLIP textual feature. This requires nothing more than a pre-trained contrastive model and a large set of sole text samples to craft a powerful captioner.
@@
In this paper, we show that many zero-shot captioners, paired with the right components, can be easily restructured to perform zero-shot region-based captioning. Therefore, we identify and study in detail the most critical components of this novel zero-shot regional captioning framework. Particularly, we focus our attention on the pre-trained vision-language contrastive backbone, which should be able, unlike CLIP, to create meaningful patch representations. To this aim, we largely explore DINO-based  [7, 45] variants, having better localized capabilities than CLIP. On top of this, we address multiple modality-gap mitigation strategies helping the text decoder to correctly interpret visual features without the need for paired image-text data, as well as a study on different patch aggregation methods.
@@
By analyzing existing components and employing vision backbones able to output patch-level meaningful representations like DINO, we show that we can enable many zero-shot captioners to reach state-of-the-art or comparable results in many zero-shot captioning task variants requiring captioning sub-parts of the entire image — dense captioning [23], region-set captioning [10], up to the standard image captioning [59] where the region to caption extends over the entire image. To better showcase the effectiveness of our framework in extreme patch-based captioning scenarios, we also introduce the trace captioning task, requiring the captioning of an image region specified by a mouse trace.
@@
To summarize, our contributions are the following: a) we reformulate captioning by shifting perspective from the image-to-caption approach to a patch-to-caption one, unifying local and global tasks in one framework which does not require region-level supervision, through the exploitation of frozen vision-language backbones, b) we repurpose existing models to work within this novel framework, by analyzing the role of the key components, with a special attention to the vision backbone, c) we show the performance of these models on four zero-shot captioning tasks spanning different region granularity, from captioning few patches to the whole image, showing the effectiveness of the proposed perspective shift proposed by our framework despite its simplicity.
@@
Figure 1: Patch-centric framework for unified zero-shot captioning. A. Overview of our framework. First, we extract language-aligned dense patch embeddings from the image using a VLM. Given a region, we select the underlying patches and aggregate their features to obtain a region representation. Finally, we obtain the region caption by applying a zero-shot text decoder, that is a) conditioned on the latent region representation, b) trained on text-only data, and c) equipped with a mechanism to handle the modality gap present in vision-language common spaces. This enables regional captioning without requiring region-level supervision. B. By aggregating patch-level features from arbitrary image regions, we can flexibly handle multiple captioning tasks across spatial granularities in a unique model.
@@ 2 Related Work
Language-aligned Dense Image Representations are crucial for our goal of captioning at patch level. Vision-language models (VLM) like CLIP [50] introduced a powerful approach to learning global modality representations in a shared space via contrastive learning, paving the way to solve several downstream tasks, including captioning [40, 11]. However, in zero-shot settings, CLIP-like representations are known to struggle with dense tasks due to misalignment between local visual patches and fine-grained semantics [71, 52, 6]. On the other hand, visual-only self-supervised models (SSM) like DINO [7, 45] excel in local semantic modeling but lack a bridge with language. Recent works like SILC [42], DINO.txt [24] and SigLIP 2 [60] attempt to get the best of both worlds by combining DINO- and CLIP-like training objectives, aiming to obtain language-aligned dense representations. INViTE [8] modifies CLIP’s visual encoder by zeroing the attention weights from each patch to all the others in the last layers, leading to more semantic patch features. DenseCLIP [53] and RegionCLIP [71] extend CLIP with additional region-level supervision. RegionCLIP leverages a region-proposal network to construct region-text pairs from image-text datasets, while DenseCLIP introduces a pixel-to-text matching loss to strengthen the alignment between local regions and textual concepts. Other methods instead exploit already existing VLMs and SSMs to get the same properties with minimal or no training: Talk2DINO [5] connects language to the DINOv2 space by mapping CLIP textual representations to DINOv2 patches. ProxyCLIP [29] instead leverages DINO’s attention maps to improve the local properties of the CLIP visual embeddings of patches.
@@
Zero-shot Image Captioning methods rely mostly on global CLIP representations to guide text generation. Early-guided decoding methods take CLIP visual features as input and introduce adaptation techniques to reduce the visual-textual modality gap [36]. DeCap [33] projects CLIP visual features into a more text-aligned space using a memory of texts as basis, while CapDec [43] and CLOSE [18] inject noise during text-only training to enable decoding also from the CLIP visual space. Diffusion Bridge [30] further mitigates this gap by applying a diffusion model to refine visual features toward the textual manifold. To improve the generation ViECap [15], MeaCap [68], MERCap [67] and EntroCap [63] leverage external knowledge to condition the decoding together with the CLIP image representation. Late-guided decoding methods instead use CLIP as a scoring or optimization signal rather than direct input. ZeroCap [59] leverages CLIP gradients to steer the cached context during text generation, while MAGIC [56] optimizes token selection based on CLIP similarity scores. However, all the above approaches rely on global representations, which are not well-suited for capturing localized semantic details, making them suboptimal for patch-level or region-level captioning in zero-shot settings.
@@
Region-level Captioning comprises several tasks in which models are asked to produce natural language descriptions based on sub-parts of an image. They pose additional challenges as naively captioning the cropped regions or feature maps often induces a loss of the global context of the image and, thus, misinterpretation of the region. For this reason, zero-shot solutions to this family of problems are still underexplored. For controllable captioning [10] — the generation of an image caption controlled by a set or sequence of regions — and dense captioning [23] — the localization and captioning of salient regions of an image — state-of-the-art solutions like CAG-Net [65], GRiT [62], ControlCap [70], and FlexCap [14] provide good performance but need supervision with ground-truth boxes. Recent works [19], [21] moved towards the direction of arbitrary regions captioning exploiting region-level supervision, and the Localized Narratives dataset [48] — comprising images, timed captions, and timed mouse tracks — provide the ingredients for evaluating captioning also at track- or patch-level. We propose a unique framework to tackle captioning at various granularities, from image- to patch-level, in a zero-shot setting.
@@ 3 A Patch-centric Framework
Overview.
Figure 1 provides an overview of our framework, which reformulates captioning around image patches as the fundamental visual units. An input image is first encoded by a frozen vision–language backbone into dense, language-aligned patch embeddings. A region is then represented by aggregating the embeddings of its constituent patches, and a text decoder generates a natural-language caption from this aggregated representation.
Assumptions.
Our approach operates under three main assumptions: (i) no region–text supervision is available during training, meaning the model never observes captions paired with localized image regions, (ii) the vision–language backbone remains frozen, providing patch-level representations that reside in, or can be projected into, a joint embedding space shared with the text encoder, and (iii) no image-text supervision is available to train the text decoder, which instead can be trained exclusively on text data, learning to reconstruct captions from their textual embeddings.
At inference, the decoder interprets visual features through the shared multimodal space, enabling region-level captioning without any region–text or image-text annotations. The following subsections detail how each component is instantiated, how regions are represented through a parameter-free patch aggregation, and how we address the modality gap between visual and textual representations.
@@ 3.1 Motivation and Formulation
Traditional regional captioning [23, 14, 70] follows an early injection of region specification in the model. Formally, given an image $I$ and a region $R$, the region caption $t$ is modeled as $t=\mathcal{D}(I,R)$. Such models require region-caption annotations and often use dedicated models or losses per captioning task or granularity. A step forward can be moved by introducing a formulation which disentangles image encoding and postpones region specification, defining $t=\mathcal{D}(\psi(I),R)$, where $\psi(I)$ provides a visual representation of the image independent of the region $R$, and $\mathcal{D}$ performs late region selection and text decoding. In order to avoid training the whole pipeline using region-level labels, we further decompose the decoder module $\mathcal{D}(\cdot)$ into two distinct modules: a parameter-free fixed aggregation $\operatorname{agg}_{R}$ of patch-level representations, and an actual text decoder $\phi(\cdot)$ not directly conditioned on regions $R$, so that $t=\phi(\operatorname{agg}_{R}(\psi(I),R))$. We will detail these two factorized components in the following paragraphs.
@@
We train a text decoder $\phi:\mathbb{R}^{D}\to\mathcal{T}$ using a prefix language modeling approach, where the decoder reconstructs a caption $t\in\mathcal{T}$ from its text embedding $\psi_{\text{t}}(t)$. If $\psi_{\text{t}}(t)$ is aligned to the visual encoder $\psi_{v}(I)$, we could directly decode the visual embeddings using a text-only decoder trained on text-only data. In such a case, we can finally decode the region representation into a natural language caption $t=\phi(\mathbf{v}_{R})$. However, the assumption that $\phi(\cdot)$ can digest features from $\psi_{v}(I)$ while being trained to reconstruct features from $\psi_{\text{t}}(t)$ is often too optimistic due to the prominent modality gap [36]. In fact, text and image representations, despite being in the same multimodal space, occupy different, separated subspaces. To obtain a decoder applicable to visual embeddings, we experimented with three mitigation strategies for this gap. The first strategy, following [33], introduces a projection step at inference that maps visual features into the text subspace using a memory of text embeddings. The second, inspired by [18] and [43], trains the decoder under input perturbations, enforcing robustness to visual embeddings laying in another subspace. Finally, following [30], we explore a diffusion-based strategy that applies a diffusion model trained on the textual manifold to the patch-level representation, effectively "bridging" the gap by iteratively refining the visual features toward the text subspace. We analyze the effects of the modality gap mitigation strategies in SM§12.
Overall, our formulation offers three main advantages: a) it is zero-shot by design, in the sense that it only requires textual descriptions to be trained and does not require paired image-text samples to train the text decoder, b) any region (e.g., whole image, box, mask, free-form trace, single point) is addressed identically, supporting a modular, general-purpose regional captioning pipeline, and c) our method requires only a single forward pass of the vision backbone to extract patch features for an entire image, which can then be reused to caption multiple regions without rerunning the full pipeline for each one.